OpenStudy (anonymous):

F^-1(y) F(x)= x/6+ 3

OpenStudy (mathmale):

Hint: I'd be a lot more willing to address this question if only you'd include the directions.

OpenStudy (anonymous):

y=x/6+3 y-3=x/6 x=6y-18 so F(x)-1=6x-18

OpenStudy (mathmale):

Please, include verbal explanations of what you're doing. The original problem posted lacks directions and your "solution" lacks explanations (as well as a conclusion).

OpenStudy (anonymous):

@youngstar10 understand?